### 15 Dec 2023 v1

# RDR: the Recap, Deliberate, and Respond Method for Enhanced Language Understanding

Abstract
Natural language understanding (NLU) using neural network pipelines often requires additional context that is not solely present in the input data. Through Prior research, it has been evident that NLU benchmarks are susceptible to manipulation by neural models, wherein these models exploit statistical artifacts within the encoded external knowledge to artificially inflate performance metrics for downstream tasks. Our proposed approach, known as the Recap, Deliberate, and Respond (RDR) paradigm, addresses this issue by incorporating three distinct objectives within the neural network pipeline. Firstly, the Recap objective involves paraphrasing the input text using a paraphrasing model in order to summarize and encapsulate its essence. Secondly, the Deliberation objective entails encoding external graph information related to entities mentioned in the input text, utilizing a graph embedding model. Finally, the Respond objective employs a classification head model that utilizes representations from the Recap and Deliberation modules to generate the final prediction. By cascading these three models and minimizing a combined loss, we mitigate the potential for gaming the benchmark and establish a robust method for capturing the underlying semantic patterns, thus enabling accurate predictions. To evaluate the effectiveness of the RDR method, we conduct tests on multiple GLUE benchmark tasks. Our results demonstrate improved performance compared to competitive baselines, with an enhancement of up to 2% on standard metrics. Furthermore, we analyze the observed evidence for semantic understanding exhibited by RDR models, emphasizing their ability to avoid gaming the benchmark and instead accurately capture the true underlying semantic patterns.
**footnotetext: These authors contributed equally to this work.
Introduction
Previous research in the field of natural language understanding (NLU) and neural network pipelines has acknowledged the necessity of incorporating additional context beyond input data (Sheth et al. 2021). Studies have recognized that achieving accurate language understanding and interpretation requires more than just input data. To address this limitation, one well-established approach is to integrate external knowledge graphs as supplementary context (Zhu et al. 2023). These knowledge graphs contain structured information about entities, relationships, and concepts. This enables the neural network to infer semantic connections between entities, even when such relations are not explicitly stated in the data alone, thus uncovering implicit or missing contextual associations.
However, a notable concern raised by previous research pertains to the vulnerability of NLU benchmarks to manipulation by neural models (Bender and Koller 2020; McCoy, Pavlick, and Linzen 2019). This issue casts doubt on the reliability and generalizability of the performance metrics reported by neural models on benchmark datasets and has undergone extensive scrutiny in the NLU community. Researchers have explored various methods to mitigate benchmark gaming and ensure that NLU models exhibit authentic language understanding. These methods include introducing more comprehensive evaluation protocols, employing adversarial testing, and designing different types of knowledge-infusion architectures. However, unified evaluation of the integrity of these diverse methodologies remains a challenge.
We propose a novel approach called the Recap, Deliberate, and Respond (RDR) paradigm, which addresses these limitations by integrating three distinct objectives within the neural network pipeline. The first objective, Recap, involves paraphrasing the input text using a dedicated model. This process captures the essential information and essence of the input. The second objective, Deliberation, focuses on encoding external graph information related to entities mentioned in the input text. This step utilizes a graph embedding model to leverage the knowledge within external knowledge graphs. By integrating this external context into the neural network pipeline, the Deliberation objective enhances the model’s ability to comprehend the relationships between entities and extract relevant information for downstream tasks. The final objective in the RDR paradigm is Respond, which employs a classification head model. This model utilizes representations from the Recap and Deliberation modules to generate the final prediction. By incorporating insights gained from the Recap and Deliberation stages, the Respond objective enables more accurate and informed predictions. The cascading structure of the three objectives, combined with the minimization of a combined loss, prevents models from exploiting statistical artifacts to artificially inflate performance metrics. This robust methodology facilitates capturing the true underlying semantic patterns within the combined data and external knowledge, leading to more reliable and accurate predictions.
Figure 1: In the figure (top), a tradiational neural network pipeline is depicted, enhanced with external knowledge to handle GLUE tasks such as entailment, similarity, and other types of natural language inference (e.g., contradiction and neutral). Initially, the tokenized text undergoes encoding by a language model, resulting in an embedding. Following that, a method based on graph embedding is employed to extract and embed a subgraph that is relevant to the input text. This involves extracting entities within a certain distance threshold from the entities present in the text. Subsequently, the two embeddings—the language model embedding and the graph embedding—are merged and passed through a classification head to obtain the predicted logits. To train this model, the cross-entropy loss between the predicted logits and the actual output is minimized. In the RDR paradigm (bottom), the tokenized input also goes through a paraphrasing model, and a paraphrasing loss is calculated. Additionally, the graph-embedding-based subgraph extraction method is compared to a ground truth subgraph, and a graph embedding loss is computed. The total loss is the sum of the losses from the paraphrasing loss, graph embedding loss, and classification head loss.
To evaluate the effectiveness of the RDR paradigm, tests were conducted on multiple GLUE benchmark tasks relating to sentence similarity, textual entailment, and natural language inference (Wang et al. 2018; Sharma et al. 2019; Demszky, Guu, and Liang 2018; Dolan and Brockett 2005; Poliak 2020). The results demonstrated superior performance compared to competitive baselines, with improvements of up to 2% on standard evaluation metrics. These findings highlight the capability of the RDR approach to enhance NLU performance. Furthermore, we discuss, using some example inferences, the capacity of models trained using the RDR methodology for semantic understanding that is robust to statistical artifacts. We find that the introduction of the Recap and Deliberation objectives leads to better comprehension of the underlying semantic patterns in the combined data and external knowledge for the example inferences. Figure 1 shows a comparison of a traditional training pipeline for the previously mentioned GLUE tasks vs. the RDR method.
RDR-Methodology
Traditional Method
Notations: Functions, and their Inputs and Outputs
  • Input Text: $x$, Tokenized Text: $T(x)$
  • Language Model, Input, Function and Output: $x^{\prime}=f(T(x),\theta)$
  • Subgraph Extractor, Input, Function, and Output:
    $KG_{x}=subgraph\_extract(T(x),KG)$
    , here $KG$ is a large knowledge graph (e.g., ConceptNet).
  • Graph Embedding Model, Input, Function and Output: $e_{x}=Aggr(g(KG_{x},i\in KG_{x},\theta^{\prime}))$, here $Aggr$ is an aggregation function (e.g., average of all node embeddings in $KG_{x}$, and $i\in KG_{x}$ denotes the nodes in the subgraph $KG_{x}$.
  • Embedding Fusion Model with Classification Head, Input, Function, and Output: $z=h(e_{x},x^{\prime},\theta^{{}^{\prime\prime}})$
  • Loss: Cross Entropy (CE) loss with ground truth denoted as $y$, $CE(z,y)$.
Forward Pass and Loss Calculation During Training
The tokenized text $T(x)$ is fed into a language model, resulting in the embedding $x^{\prime}$.
An off-the-shelf graph extraction method is applied to extract a subgraph $KG_{x}$ from the larger knowledge graph $KG$, and, subsequently, a graph embedding model $Aggr(g(KG_{x},\theta^{\prime}))$ is used to obtain the graph embedding for $x$, denoted as $e_{x}$. The language model embedding $x^{\prime}$ and the subgraph embedding $e_{x}$ are passed through an embedding fusion model with a classification head $h(e_{x},x^{\prime},\theta^{{}^{\prime\prime}})$ to obtain the logits $z$. The loss is computed using the logits $z$ and the ground truth $y$.
RDR Method
Notations: Functions, and their Inputs and Outputs
  • Input Text: $x$, Tokenized Text: $T(x)$
  • Paraphrasing Model, Input, Function and Output: $x^{\prime}=\\ f(T(x),\theta)$
  • Paraphrasing Loss: The discrepancy between the paraphrased text $x^{\prime}$ and the original text $x$ measured using cross entropy loss between the logits from the model $f$ and the ground truth distribution of tokens in $x$. We denote this loss as $PL(x^{\prime},x)$.
  • Subgraph Extractor, Input, Function, and Output:
    $KG_{x}=subgraph\_extract(T(x),KG)$
    , here $KG$ is a large knowledge graph (e.g., ConceptNet).
  • Graph Embedding Model, Input, Function and Output: $e_{x}=Aggr(g(KG_{x},i\in KG_{x},\theta^{\prime}))$, here $Aggr$ is an aggregation function (e.g., average of all node embeddings in $KG_{x}$, and $i\in KG_{x}$ denotes the nodes in subgraph $KG_{x}$.
  • Embedding Loss Calculator: First, all links in $KG_{x}$ are predicted using the model $Aggr(g(KG_{x},i\in KG_{x},\theta^{\prime}))$, where the presence of a link between two nodes $i$ and $j$ is determined by if $||g(KG_{x},i,\theta^{\prime})-g(KG_{x},j,\theta^{\prime})||\leq\tau$, where $tau$ is a “closeness” threshold set empirically. Then link prediction metrics are calculated, e.g., mean reciprocal rank (MRR), hits@k, etc. This loss is denoted as $GEL(x,KG_{x})$
  • Embedding Fusion Model with Classification Head, Input, Function, and Output: $z=h(e_{x},x^{\prime},\theta^{{}^{\prime\prime}})$
  • Response Loss: Cross Entropy (CE) loss with ground truth denoted as $y$, $RL(z,y)$.
  • Total Loss: $L=PL(x^{\prime},x)+GEL(x,KG_{x})+RL(z,y)$.
    Forward Pass and Loss Calculation During Training
    The tokenized text $T(x)$ is fed into a language model, resulting in the embedding $x^{\prime}$. Next, the paraphrasing loss $PL(x^{\prime},x)$ is calculated. An off-the-shelf graph extraction method is applied to extract a subgraph $KG_{x}$ from the larger knowledge graph $KG$, and, subsequently, a graph embedding model $Aggr(g(KG_{x},\theta^{\prime}))$ is used to obtain the graph embedding for $x$, denoted as $e_{x}$. Also, the link prediction loss $GEL(x,KG_{x})$ is calculated. The language model embedding $x^{\prime}$ and the subgraph embedding $e_{x}$ are passed through an embedding fusion model with a classification head $h(e_{x},x^{\prime},\theta^{{}^{\prime\prime}})$ to obtain the logits $z$. The response loss is computed as $RL(z,y)$. Finally, the total loss is computed as $L=PL(x^{\prime},x)+GEL(x,KG_{x})+RL(z,y)$.
Experiments and Results
Figure 2: Illustration of the process of extracting subgraphs from the knowledge graph given an input instance. For the pre-trained Graph Encoder Network, we use ConceptNet’s Numberbatch embeddings and use a span length of three in our experiments.
Our implementation utilized task-specific hyperparameters, previously identified as optimal for the GLUE benchmark, except for training until the completion of 1 Epoch. Throughout the training and evaluation process, a batch size of 8 was employed. Additionally, we integrated task-specific knowledge graphs or subgraph representations, amounting to up to 10% of the data. Our approach adhered to the predefined train-validation split established by the GLUE benchmark. The reported results represent the average of two independent runs. The knowledge graph utilized in this study encompassed DBPedia, ConceptNet, Wiktionary, WordNet, and the OpenCyc Ontology. These knowledge graphs consist of interconnected objects and their relationships, forming semantic associations (Auer et al. 2007; Speer, Chin, and Havasi 2017; Matuszek et al. 2006). Figure 2 illustrates the process of extracting subgraphs from the input text.
Approximately 300K subgraphs were obtained from the knowledge graphs across all inputs. The relationships discovered include Antonym, DistinctFrom, EtymologicallyRelatedTo, LocatedNear, RelatedTo, SimilarTo, Synonym, AtLocation, CapableOf, Causes, CausesDesire, CreatedBy, DefinedAs, DerivedFrom, Desires, Entails, ExternalURL, FormOf, HasA, HasContext, HasFirstSubevent, HasLastSubevent, HasPrerequisite, HasProperty, InstanceOf, IsA, MadeOf, MannerOf, MotivatedByGoal, ObstructedBy, PartOf, ReceivesAction, SenseOf, SymbolOf, and UsedFor.
We experiment with the best-performing models from huggingface with $\leq$ 500 million parameters, i.e., BERT, RoBERTa, and ALBERT. We note that RoBERTa is especially suited for entailment tasks (Devlin et al. 2018; Liu et al. 2019; Lan et al. 2019). Table 1 shows the accuracies for all the models with and without the RDR training method. For the graph embedding model in the RDR model, we use the TransE algorithm (Bordes et al. 2013). We find that the RDR method shows significant improvements over the baselines, with just 10% of the knowledge graphs used. We choose a random 10% of the knowledge graph for each training run to avoid the observed phenomenon of gaming of evaluation benchmarks by language understanding models, i.e., to avoid the model fitting spurious patterns afforded to it by the additional knowledge as a means to achieve high accuracy scores.
M MNLI-M MNLI-MM QNLI QQP WNLI MRPC RTE
B
82.44
83.52
90.49
90.1
54.92
82.11
66.06
RB
83.31
84.47
91.25
90.78
56.33
82.86
67.87
R
85.07
85.19
91.06
90.17
56.33
86.03
62.81
RR
85.78
85.95
91.85
90.96
57.75
86.76
63.53
A
84.34
85.32
90.6
90.25
57.75
86.27
66.06
RA
85.03
85.82
91.18
90.74
59.15
87
66.43
Table 1: RB: RDRB, RA: RDRA, RR: RDRR, M: MODEL, MNLI-M: MNLI-MATCHED, MNLI-MM: MNLI-MISMATCHED, B: BERT-BASE, R: ROBERTA-BASE, A: ALBERT-BASE-V2. Results for RDR method compared to models that do not use the RDR method. We see improvements of up to 2%, on average 1% using only 10% of the knowledge graphs showing promise of the RDR methodology for improved language understanding.
Conclusion and Future Work
This paper presents the formalization, and initial experimental outcomes of a new training approach known as ”Recap, Deliberate, and Respond” (RDR). We see that RDR achieves superior performance compared to baseline methods, while also displaying resilience against manipulation of benchmarks (as evidenced by observing performance improvements when using only a random 10% of the available knowledge during each training iteration). Subsequent research will explore the utilization of diverse knowledge sources, including domain-specific knowledge, broader general knowledge (e.g., from Wiki), and others. Additionally, future work will involve analyzing the augmenting ability of using SOTA large-scale models & diverse geometrical embeddings within the RDR framework.
Acknowledgements
This research is built upon prior work (Zi et al. 2023; Rawte et al. 2022; Roy et al. 2023c; Venkataramanan et al. 2023; Roy et al. 2023a, 2022, a), and supported in part by NSF Award 2335967 ”EAGER: Knowledge-guided neurosymbolic AI” with guardrails for safe virtual health assistants”. Opinions are those of the authors and do not reflect the opinions of the sponsor (Roy et al. 2021, 2023b; Sheth et al. 2021, 2022; Sheth, Roy, and Gaur 2023).